What are the intercepts of the equation? Graph the equation.
-3x + 4y = -12

A. x-intercept: (4, 0)
y-intercept: (0, –3)


B. x-intercept: (–3, 0)
y-intercept: (0, 4)


C. x-intercept: (–3, 0)
y-intercept: (0, 4)

D. x-intercept: (4, 0)
y-intercept: (0, –3)




Answer by waynest(281)

-3x + 4y = -12
make y = 0
-3x + 4(0) = -12
-3x + 0 = -12
-3x = -12
-3x/3 = -12/3
x = 4
(4,0)
make x = 0
3(0) + 4y = -12
4y = -12
4y/4 = -12/4
y = -3
(0,-3)
therfore A or D are correct